Contribution Margin and Contribution Margin Ratio

For a recent year, McDonald’s (MCD) company-owned restaurants had the following sales and expenses (in millions):

Sales $38,300
Food and packaging $(10,274)
Payroll (9,700)
Occupancy (rent, depreciation, etc.) (11,576)
General, selling, and administrative expenses (5,600)
$(37,150)
Operating income $1,150

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What is McDonald's contribution margin? Round to the nearest million. (Give answer in millions of dollars.)
$ million

b. What is McDonald's contribution margin ratio?
%

c. How much would operating income increase if same-store sales increased by $2,300 million for the coming year, with no change in the contribution margin ratio or fixed costs? Round your answer to the closest million.
$ million
Answer #1

Variable costs = Food and packing + Payroll + (40%* general selling and administrative expenses)

= 10274+9700+(40%*5600) = 22214

Contribution margin = Sales - Variable costs

= 38300 - 22214

= 16086

Contribution margin ratio = Contribution margin/Sales

= 16086/38300

= 42%

Increase in operating income = 2300 million * 42%

= 966 million